colors.tex

Bent Petersen

Oct 24, 2000

The easiest way to use color text in L^{A}TEX 2ε is to use the color package. You load the color package in your preamble as usual

\documentclass[11pt]{article}
\usepackage[usenames]{color}

Here the option [usenames] causes the definitions for the 68 colors known to dvips (see next page) to be preloaded. Those colors can then be specified by name. For example {\color{BrickRed} BrickRed} produces BrickRed. Alternately one can use the command

\textcolor{BrickRed}{this is BrickRed}

which produces this is BrickRed. If the 68 predefined colors do not suffice, or if you do not want to load the definitions of all of them, you may define your own custom colors. Here is a dark blue defined in terms of the RGB (red, green, blue) color model.

\definecolor{MyDarkBlue}{rgb}{0,0.08,0.45}
\begin{center}
{\color{MyDarkBlue}This color is MyDarkBlue}
\end{center}

This color is MyDarkBlue Once we have defined the new color we can use it anywhere in the current

document. For example, \textcolor{MyDarkBlue}{a nice blue} produces

a nice blue. Custom colors may also be defined in terms of the CMYK (cyan, magenta, yellow, black) color model. Here is a light magenta.

\definecolor{MyLightMagenta}{cmyk}{0.1,0.8,0,0.1}
\begin{center}
{\color{MyLightMagenta}This color is MyLightMagenta}
\end{center}

This color is MyLightMagenta

The 68 standard colors known to dvips

Apricot  Aquamarine  Bittersweet Black
Blue  BlueGreen  BlueViolet BrickRed
Brown  BurntOrange CadetBlue CarnationPink
Cerulean  CornflowerBlue Cyan  Dandelion
DarkOrchid  Emerald  ForestGreen Fuchsia
Goldenrod  Gray  Green  GreenYellow
JungleGreen  Lavender  LimeGreen Magenta
Mahogany  Maroon  Melon  MidnightBlue
Mulberry  NavyBlue  OliveGreen Orange
OrangeRed  Orchid  Peach  Periwinkle
PineGreen  Plum  ProcessBlue Purple
RawSienna  Red  RedOrange RedViolet
Rhodamine  RoyalBlue  RoyalPurple RubineRed
Salmon  SeaGreen  Sepia  SkyBlue
SpringGreen  Tan  TealBlue  Thistle
Turquoise  Violet  VioletRed  White
WildStrawberry Yellow  YellowGreen YellowOrange